BeautifulAngel:

5x2 + 5y2 − 20x + 30y + 40 = 0 How do I get the simplified form? ( r = ? )

BeautifulAngel:

@Hero

Hero:

Looks like you need to convert that to the form \(x^2 + y^2 = r^2\). First start by dividing both sides by 5. Let me know what you get.

BeautifulAngel:

How do I do that? Math is my worst subject..

Hero:

Simply divide each term by 5 on both sides.

BeautifulAngel:

Both sides? What do you mean both sides? Please give an example??

Hero:

What I mean is, Divide by 5 on Both sides of the EQUAL sign.

Hero:

On the left side of the equation remember that \(\dfrac{a + b}{c} = \dfrac{a}{b} + \dfrac{b}{c}\) For example, suppose you have \(6a + 3b\) and need to divide both terms by 3. Perform the following steps: \(\dfrac{6a + 3b}{3} = \dfrac{6a}{3} + \dfrac{3b}{3}\) Doing so results in \(2a + b\)

Hero:

@BeautifulAngel Still lost?

Hero:

Okay, now that you're back can we please finish?

BeautifulAngel:

Yes, of course. Let me do this dividing real quick

BeautifulAngel:

x^2+y^2-4x+6y+8=0 ???

Hero:

Correct so far.

Hero:

Now you just have to do that trick I showed you earlier.

BeautifulAngel:

What trick?

Hero:

See your previous question

BeautifulAngel:

I looked, but I can't figure out how to set it up.. >.< ugggh!!!!!!!!!!!!! I'm so freaking dumb, I'm going to fail this class

BeautifulAngel:

Do I divide -4 and 6 by 2 and replace x with the answer?

Hero:

Oh, sorry. I hope you're still here. I had to deal with something very important.